Tom's father is three times as old as Tom. Four years ago Tom's father was four times as old as Tom. How old is Tom?
Answer by AnlytcPhil(1116)

Tom's father is three times as old as Tom. Four years ago
Tom's father was four times as old as Tom. How old is Tom?


Let:
Tom's age now = T
>>...Tom's father is three times as old as Tom...<<
TRANSLATION:
Tom's father's age now = 3T
Then
Tom's age four years ago = T-4
and
Tom's father's age four years ago = 3T-4
>>... Four years ago Tom's father was four times as old as Tom...<<
TRANSLATION:
Tom's father's age four years ago = 4 times Tom's age 4 years ago
3T-4 = 4(T-4)
3T-4 = 4T-16
3T-4T = 4-16
-T = -12
T = 12
Check:
Tom is 12 years old. His father is 3 times as old as Tom, or 36.
Four years ago, Tom was 8 and his father was 32. So indeed Tom's father
was four times as old as Tom was then.
Edwin